find the equation of tangents from the point 2,2 to the ellipse 4x2+9y2=36

Grade:11

1 Answers

bharat bajaj IIT Delhi
9 years ago
In General, if the ellipse is :
x2/a2+y2/b2= 1
The equation of tangent is :
(xx1)/a2+(yy1)/b2= 1.
Hence, for this ellipse, the equation of tangent is :
8x + 18y = 36
or 4x + 9y = 18
Thanks & Regards
Bharat Bajaj
askiitian faculty
IIT Delhi